OpenStudy (anonymous):

After I used quadratic formula for -3x^2+6x-5

OpenStudy (campbell_st):

well I'd use the discriminant \[b^2 - 4ac \] because if \[b^2 - 4ac < 0 \] and a is negative, then the curve is negative definite so 36 - 4 x (-3)(-5) < 0 -24 < 0 so the curve has no real roots... and since a = -3 the curve is concave down. the curve has an axis is symmetry at x = 1 and a vertex at (1, -2) so the curve looks like  so \[-3x^2 + 6x - 5 \le 0, ..(-\infty, \infty)\]

OpenStudy (campbell_st):

or the curve is <= 0 for all real x

OpenStudy (anonymous):

wow...thanks! why did you use only b^2 - 4ac?

OpenStudy (campbell_st):

because the discriminant gives you information about the parabola > 0 means real roots which can be rational or irrational = 0 means equal or repeated roots. < 0 no real roots. so once you know there are no real roots... the curve is always positive ( positive definite) or in your case always negative ( negative definite) actually my 1st thought was to graph the curve, when I saw you results for the GQF... I knew it was always negative... but just needed to show you

OpenStudy (anonymous):

how did you get symmetry at x = 1 and a vertex at (1, -2) ?

OpenStudy (campbell_st):

ok... the line of symmetry for a parabola, y = f(x) is found using \[x =\frac{-b}{2a}\] which you may have seen in the GQF... so x = -6/(2x-3) x = 1

OpenStudy (anonymous):

so. much. detail. hehe. when finding the answer, would it be like x<or equal to something? i'm a bit confused.

OpenStudy (campbell_st):

ok.... lets look at a couple x\[x^2 + 5x - 6 < 0\] so \[(x + 6)(x - 1) < 0\] graphing it  so the question is asking when is the curve below the x axis -6 < x< 1 is the algebraic solution. in these questions you are only seeking solutions for real numbers... if you get a complex solution, then the curve doesn't cut the x- axis